I need help with this exercise.

Two countries, Canada and Indonesia, produce two goods, Computers (C) and rice (R) with labour (L) as the only input. The quantities of input required to produce a unit of each output are as follows:

Computers

Rice

Indonesia

8

1

Canada

10

2

L in Canada is = 30, in Indonesia is =40.

i) Sketch the PPF in each country (with QC on the horizontal axis). Indicate the intercepts and slope of the PFF for each country.

ii)What are the relative prices in each country?

The two countries open to international trade. The world relative price of computers is 7.

iii) What is the pattern of specialization? Of trade?

Thanks to technological progress, the productivity of Indonesia in producing computers increases. Now the input requirement in the sector is equal to 4.

iv) Assume a new equilibrium value of world relative price. Will Canada gain by returning to autarky?
Answer #1

Maximum number of quantities produced(total available labor divided by labour required to produce one unit)

  1. Computers

    Rice

    Indonesia

    40/8=5

    40/1=40

    Canada

    30/10=3

    30/2=15

i ) answer:

ii. answer:

relative price in Indonesia

price of computer x maximum possible quantity of computer= price of rice x maximum possible quantity of rice

hence, the price of rice/price of computer= maximum possible quantity of computer/maximum possible quantity of rice=5/40=1/8

and the price of computer/price of rice=8

relative price in Canada

price of computer x maximum possible quantity of computer= price of rice x maximum possible quantity of rice

hence, the price of rice/price of computer= maximum possible quantity of computer/maximum possible quantity of rice=3/15=1/5

and the price of computer/price of rice=5/1

iii) answer:

if the relative price of computer=7

then Canada will have a comparative advantage here. because they can produce computer at a lower relative price. hence they will specialise at producing computers.

iv).answer:

now technology changed and Indonesia now requires only 4 units of labour to produce one unit of a computer. now the production table looks like

Maximum number of quantities produced(total available labour divided by labour required to produce one unit)

Computers

Rice

Indonesia

40/4=10

40/1=40

Canada

30/10=3

30/2=15

now relative price of computer at indonesia= 40/10=4

canada's relative computer at indonesia=15/3=5

now a probable world price for computer is 4.5

even now canada doesn't gain by going back to autarky. they'll be better off by specialising in rice production.